A rectangular lawn is 45 feet long and 30 yards wide. Find the perimeter in feet. Then find the perimeter in yards.
Answer by checkley77(12844)

2*45+2*30*3=90+180=270 feet is the perimeter.
2*45/3+2*30=2*15+60=30=60=90 yards is the perimeter.